Shift summary for day 7/11/99

Jason Hamm (jhamm@physics.arizona.edu)
Sun, 11 Jul 1999 13:30:45 -0700 (MST)


Shift Summary for day shift, 07/11/99

On Shift: J. Hirschauer, J. Hamm

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: DA problem
========================================================================
Runs ended this shift:

Run 13856 - BEAM_832 (18.3 M events)

Tapes brought to FCC this shift: QK*113-115, QKB039
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: 832 data taking at ~6.5 E12
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13856 7:35 DDD3 veto enabled. Gateway still running,
can't clear busy. Ed advises us to start new run.
Stop process hangs - reason still unknown. Tapes
almost full anyways, so fresh start and load
a new tape set.

8:03 Start beam832 run 13857

13857 9:25 Laser crate DYC busy. Kfrend twice - now we get
pipeline errors. Quick pipeline reset fixes it.

13857 9:50 MU2 DYC busy. Kfrend - now DC4/SA4 busy.
Kfrending again fixes it.

13857 10:05 Onsplit requests QKCC05 to be mounted on Ktev2.
We obey.

13857 11:15 SDAQ shows +5V supply for UCLA shaper-splitter for
HA is out of range at +5.37V. Message disappears
after about 10 minutes. SET is informed.

13857 13:05 SDAQ shows +/- 24V supply for NIM bin #2 in RR160
(trigger rack) is out of range at +/- 26V.
Everything is functioning properly, but this
should be monitored to make sure it doesn't drift
further out of range.

13857 14:44 Onsplit requests QKN010 to be mounted on Ktev4.
We do so.
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Keep an eye on the +/- 24V supply to RR160.